# Spatio-temporal Fourier Transformer (StFT) for Long-term Dynamics Prediction

### v1

@@ Appendix C Experimental Details
Training/validation/test data sets. For the plasma MHD data, we split the trajectory of $1221$ snapshots into a training set (the first $927$ snapshots), a validation set (the middle $50$ snapshots), and a test set (the last $244$ snapshots). For the Navier-Stokes dataset, we have a total of $100$ trajectories ($101$ snapshots for each trajectory), and split them into $80$ trajectories for training, $10$ for validation, and $10$ for testing. For the shallow-water dataset, we have a total of $200$ trajectories ($72$ snapshots for each trajectory), and split them into $160$ trajectories for training, $20$ for validation, and $20$ for testing.
@@
Generative residual correction block. We follow a two-step training protocol in training StFT -F: first, we train StFT thoroughly, and then we train the generative residual correction block. We employ a rectified flow to learn distributions of the residuals given the prediction of the last layer of StFT and the past snapshots. We implement a similar structure to the Diffusion Transformer (DiT) as the backbone model (Peebles & Xie, 2023). In each DiT block, we apply adaptive layer normalization before a self-attention layer and an MLP layer. We use adaLN-Zero for time conditioning. For the historical snapshots $\tilde{u}_{t}$ and the prediction $\mathcal{F}_{\theta_{d}}(\tilde{u}_{t})$ from the last layer of StFT , these conditions are incorporated as extra input tokens.
@@
Hyperparameters. For StFT on the plasma dynamics, 3D FFT is used to encode the spatio-temporal inputs in the frequency path. We use the patch size of $128$ for the the first StFT block and and $64$ for the second StFT block. The overlapping size is set to $1$. The hidden dimension is set to $128$. The depth for each StFT block is set to $6$. We keep the lowest $8$ frequencies for each spatial dimension. For the rectified flow block, the depth is set to $8$, and the hidden dimension is set to $128$. For the Navier-Stokes equation, StFT uses a patch size of $25$ for the coarse block (the first block), $13$ for the middle block, and $8$ for the last block. The overlapping size is set to 00, and the frequency path is not used. For each block in the hierarchical structure, the depth is set to 8, and we use a hidden dimension of 512. In the rectified flow block, we use a depth of $4$ and set the hidden dimension to $128$. For the shallow-water equation, three levels of StFT blocks are employed, and their patch sizes are set to $128$, $64$, and $32$ respectively. For each block, the depth is set to $6$, and the hidden dimension is set to $512$. We use the 2D FFT to encode the spatio-temporal inputs, and the lowest $8$ frequencies are kept for each spatial dimension. The overlapping size is set to $1$. For the rectified flow model, we use a depth of $8$ and a hidden dimension of $128$.
Baseline. We run enough epochs to make sure all methods are thoroughly trained. For AR-FNO, the number of modes are selected through a search in $[16,20,24]$, the number of layers are searched in $[4,5]$, and the hidden dimension is set to 256. For AR-ViT, we vary the hidden dimension from $[256,512]$, and the patch size from $[16,32,64]$. For AR-UNet, the hidden dimension of bottleneck embeddings are searched in $[64,128,256,512]$.
Training. We run enough epochs to make sure all methods are thoroughly trained. For the generative residual correction block, we run $200$ epochs after StFT is trained in the first step. We notice that when training the rectified flow on the residuals, it is crucial to normalize them, and we normalize the residuals by dividing them by their standard deviations.
Appendix D More Visualization Results
@@
Figure 8 illustrates the ground truth and predicted temporal evolution of normalized perturbed parallel vector potential $\delta A_{\parallel}$ and perturbed electron density $\delta n_{e}$ in plasam MHD using AR-StFT , AR-StFT -F, AR-FNO, AR-ViT and AR-UNet methods. StFT and StFT-F perform stable across the forecasting time horizon. Figure 9 and 10 show the pointwise error of all the models compared to the ground truth data in the Navier-Stokes and shallow-water equations, where StFT and StFT-F demonstrate lower residuals compared to other baseline models.

### v2

@@ Appendix E Experimental Details
Training/validation/test data sets. For the plasma MHD data, we split the trajectory of $1221$ snapshots into a training set (the first $927$ snapshots), a validation set (the middle $50$ snapshots), and a test set (the last $244$ snapshots). For the Navier-Stokes dataset, we have a total of $100$ trajectories ($101$ snapshots for each trajectory), and split them into $80$ trajectories for training, $10$ for validation, and $10$ for testing. For the shallow-water dataset, we have a total of $200$ trajectories ($72$ snapshots for each trajectory), and split them into $160$ trajectories for training, $20$ for validation, and $20$ for testing. To ensure a fair comparison, we impose a fixed training budget across all models. Specifically, a $48$-hour limit mesaured on one A100 GPU was set for the shallow-water equation dataset, while a $24$-hour limit is applied to both the plasma MHD and Navier-Stokes equation datasets. Table  12 summarizes the datasets and the training budget for all models.
@@
Generative residual correction block. We follow a two-step training protocol in training StFT-F: first, we train StFT thoroughly with the training budget, and then we train the generative residual correction block for another 200 epochs. We employ a rectified flow to learn distributions of the residuals given the prediction of StFT and the history snapshots. We implement a similar structure to the Diffusion Transformer (DiT) as the backbone model [44]. In each DiT block, we apply adaptive layer normalization before a self-attention layer and an MLP layer. We use adaLN-Zero for time conditioning. For the history snapshots $\tilde{u}_{t}$ and the prediction $\mathcal{F}_{\theta_{d}}(\tilde{u}_{t})$ of StFT, these conditions are incorporated as extra input tokens.
@@
Hyperparameters. For StFT on the plasma dynamics, 3D FFT is used to encode the spatio-temporal inputs in the frequency path. We use the patch size of $128$ for the the first StFT block and and $64$ for the second StFT block. The overlapping size is set to $1$. The hidden dimension is set to $128$. The depth for each StFT block is set to $6$. We keep the lowest $8$ frequencies for each spatial dimension. For the rectified flow block, the depth is set to $8$, and the hidden dimension is set to $128$. For the Navier-Stokes equation, StFT uses a patch size of $25$ for the coarse block (the first block), $13$ for the middle block, and $8$ for the last block. The overlapping size is set to 0, and the frequency path is not used. For each block in the hierarchical structure, the depth is set to 8, and we use a hidden dimension of 512. In the rectified flow block, we use a depth of $4$ and set the hidden dimension to $128$. For the shallow-water equation, three levels of StFT blocks are employed, and their patch sizes are set to $128$, $64$, and $32$ respectively. For each block, the depth is set to $6$, and the hidden dimension is set to $512$. We use the 2D FFT to encode the spatio-temporal inputs, and the lowest $8$ frequencies are kept for each spatial dimension. The overlapping size is set to $1$. For the rectified flow model, we use a depth of $8$ and a hidden dimension of $128$.
Table 13: Hyperparameter search range for each method.
Method Hyperparameter Search
DPOT Hidden dimension: $[256,512]$
Patch size: $[8,16,32]$
Depth: 6
Heads: 4
AR-FNO Modes: $[16,20,24]$
Layers: $[4,5]$
Hidden dimension: 256
AR-ViT Hidden dimension: $[256,512]$
Patch size: $[16,32,64]$
AR-UNet Bottleneck hidden dimension: $[64,128,256,512]$
Baselines. For all the baselines, we run all models with the same training time budget, as detailed in Table  12. For DPOT, we vary the hidden dimension from $[256,512]$, and the patch size from $[8,16,32]$. We set the depth as 6, and the number of heads as 4, which is the default setting from the author’s implementation. For AR-FNO, the number of modes are selected through a search in $[16,20,24]$, the number of layers are searched in $[4,5]$, and the hidden dimension is set to 256. For AR-ViT, we vary the hidden dimension from $[256,512]$, and the patch size from $[16,32,64]$. For AR-UNet, the hidden dimension of bottleneck embeddings are searched in $[64,128,256,512]$. Table  13 summarizes the hyperparameter choices for each method.
Appendix F More Visualization Results
@@
Figure 6 illustrates the ground truth and predicted temporal evolution of normalized perturbed parallel vector potential $\delta A_{\parallel}$ and perturbed electron density $\delta n_{e}$ in plasam MHD using AR-StFT, AR-StFT-F, AR-DPOT, AR-FNO, AR-ViT and AR-UNet methods. StFT and StFT-F perform stable across the forecasting time horizon. Figure 7 and 8 show the pointwise error of all the models compared to the ground truth data in the Navier-Stokes and shallow-water equations, where StFT and StFT-F demonstrate lower residuals compared to other baseline models.
Figure 11: Results of comparing StFT and StFT-F for the autoregressive prediction in $L_{2}$ relative error across different timestamps for velocity $u$ in Navier-Stokes equation. The shaded region represents the standard deviation distribution of the relative error of StFT-F. However, the uncertainty is negligible that it is not visually discernible. StFT-F demonstrates better performance in the latter stages of the forecasting time horizon compared to StFT.
In addition, we compare StFT with StFT-F for the autoregressive prediction across different timestamps for velocity $u$ in Navier-Stokes equation, shown in Figure  11. A slightly larger error is observed between $t=0.4$ and $0.6$, which can be attributed to the training objective of StFT. As StFT is optimized using a pointwise loss function, it is encouraged to produce predictions that closely match the most likely outcome. In contrast, StFT-F is designed to learn the full distribution of the target, potentially introducing higher variance in its predictions. This distributional modeling, while beneficial for uncertainty quantification, may result in marginally increased errors as it captures characteristics beyond the mean behavior.